# Similarity-Dissimilarity Loss for Multi-label Supervised Contrastive Learning

### v3

@@ Abstract
Supervised contrastive learning has achieved remarkable success by leveraging label information; however, determining positive samples in multi-label scenarios remains a critical challenge. In multi-label supervised contrastive learning (MSCL), relations among multi-label samples are not yet fully defined, leading to ambiguity in identifying positive samples and formulating contrastive loss functions to construct the representation space. To address these challenges, we: (i) first define five distinct multi-label relations in MSCL to systematically identify positive samples, (ii) introduce a novel Similarity-Dissimilarity Loss that dynamically re-weights samples through computing the similarity and dissimilarity factors between positive samples and given anchors based on multi-label relations, and (iii) further provide theoretical grounded proof for our method through rigorous mathematical analysis that supports the formulation and effectiveness of the proposed loss function. We conduct the experiments across both image and text modalities, and extend the evaluation to medical domain. The results demonstrate that our method consistently outperforms baselines in a comprehensive evaluation, confirming its effectiveness and robustness. Code is available at: https://github.com/guangminghuang/similarity-dissimilarity-loss.
Index Terms:
Multi-label Supervised contrastive learning (MSCL), multi-label classification, international classification of diseases (ICD).
I Introduction
Multi-label classification presents significant challenges due to its inherent label correlations, extreme and sparse label spaces, and long-tailed distributions. For instance, in the International Classification of Diseases (ICD) [1, 2], the presence of one label (e.g., ”Pneumococcal pneumonia”) may increase the probability of co-occurring labels (e.g., ”fever” or ”cough”). Furthermore, multi-label datasets frequently exhibit long-tailed distributions, where a small subset of labels occurs with high frequency while the majority appear rarely. This imbalance typically results in models that perform adequately on common labels but underperform on infrequent ones [3, 4]. Additionally, the number of potential label combinations increases exponentially with the number of labels, resulting in heightened computational complexity and substantial memory requirements.
@@ I Introduction
Supervised contrastive learning effectively utilizes label information to yield promising results in single-label scenarios [5]. However, identifying positive samples in multi-label supervised contrastive learning (MSCL) remains a challenge. For example, consider a set of images containing cats and puppies, wherein an anchor image depicts a cat; in the single-label paradigm, positive and negative instances can be unambiguously delineated based on their corresponding taxonomic annotations. Conversely, MSCL introduces inherent classification ambiguity when determining whether an image containing both cats and puppies should be designated as a positive or negative sample in relation to the anchor.
@@
A critical question arises: Should a sample be considered positive when its label set partially overlaps with or exactly matches that of the anchor? Currently, three principal strategies exist for identifying positive samples in multi-label scenarios [6]: (i) ALL considers only samples with an exactly matching label set as positive; (ii) ANY identifies samples with any overlapping class with the anchor as positive, and (iii) MulSupCon [6] conceptually aligns with the ANY approach but treats each label independently, thereby generating multiple distinct positive sets for individual anchor samples.
@@
However, these methods have inherent limitations, since previous research has overlooked the complicated multi-label relations among samples in MSCL. As illustrated in Figure 1, we introduce five distinct set relations among samples to facilitate a more comprehensive identification of positive sets. The ALL strategy exclusively considers relation $R2$ while disregarding the potential contributions of $R3$, $R4$ and $R5$. Furthermore, long-tailed distributions, when tail samples serve as anchors, the ALL strategy’s requirement for exact label matches significantly impedes these tail anchors from identifying adequate positive samples within a limited batch size, potentially degenerating the method to unsupervised contrastive learning in extreme scenarios [7, 8, 3]. Conversely, both ANY and MulSupCon approaches treat relations $R2$, $R3$, $R4$, and $R5$ identically with equivalent weights in contrastive loss functions, which constitutes a suboptimal approach given the inherent differences among these relations. A detailed mathematical analysis of these three methods is presented in Section II.
@@
Figure 1: Five distinct relations between samples and a given anchor. $\Omega$ denotes a universe that contains all label entities. Here is an example with five different relations between sample $p$ and anchor $i$, where the labels are represented as one-hot vectors.
@@
II Methods
In this section, we establish the preliminary notation and adhere to the conventions established in [5] to maintain consistency throughout our analysis. Subsequently, we examine the limitations of the ALL, ANY, and MulSupCon strategies and their corresponding loss functions. We then introduce our formulation of multi-label relations and present the Similarity-Dissimilarity Loss for MSCL. Furthermore, we provide a rigorous mathematical analysis to establish the theoretical foundations of the proposed methodology.
II-A Preliminaries
Given a batch of $N$ randomly sample/label pairs, $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1,\dots,N}$, where $\boldsymbol{x}_{i}$ denotes the $i$-th sample and $\boldsymbol{y}_{i}$ its corresponding labels. Here, $\boldsymbol{y}_{i}=\{y_{i}^{(l)}\}_{l=1,\dots,L}$ represents the multi-labels of sample $i$, where $y_{i}^{(l)}$ denotes the $l$-th label of sample $i$ and $L$ is the total number of labels for sample $i$. After data augmentation, the training batch consists of $2N$ pairs, $\{\tilde{\boldsymbol{x}}_{j},\tilde{\boldsymbol{y}}_{j}\}_{j=1,\dots,2N}$, where $\tilde{\boldsymbol{x}}_{2i}$ and $\tilde{\boldsymbol{x}}_{2i-1}$ are two random augmentations of $\boldsymbol{x}_{i}$ ($i=1,\dots,N$) and $\tilde{\boldsymbol{y}}_{2i-1}=\tilde{\boldsymbol{y}}_{2i}=\boldsymbol{y}_{i}$. For brevity, we refer to this collection of $2N$ augmented samples as a ”batch” [5].
II-B Multi-label Supervised Contrastive Loss
@@
Zhang et la [6] propose an approach that considers each label $\tilde{y}_{i}^{(l)}$ independently, forming multiple positive sets for a given anchor sample $i$. For each label $\tilde{y}_{i}^{(l)}\in\tilde{\boldsymbol{y}}_{i}$, the positive set for the MulSupCon is defined as:
@@
$\mathcal{L}_{i}^{mul}=\sum_{\tilde{y}_{p}^{(l)}\in\tilde{\boldsymbol{y}}_{i}}\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (6)
@@
$\mathcal{L}^{mul}=\frac{1}{\sum_{i}|\tilde{\boldsymbol{y}}_{i}|}\sum_{i\in I}\mathcal{L}_{i}^{mul}$ (7)
II-C Multi-label Relations
@@
For this example (see Figure 1), the MulSupCon successfully discriminates $R3$ and $R4$ from $R2$ and $R5$; however, it fails to establish a distinction between $R2$ and $R5$. This limitation arises primarily because MulSupCon exclusively considers the overlapping regions (Similarity  [^3] The definition of Similarity is introduced in Section II-D) between anchor $i$ and sample $p$ (i.e., The intersection of sets $\mathcal{S}$ and $\mathcal{T}$), while disregarding the complementary non-intersecting domains (Dissimilarity  [^4] The definition of Dissimilarity is introduced in Section II-D). That is to say, the similarity between positive samples and anchors is considers, but not yet dissimilarity, which is one of critical information for representation learning in MSCL.
@@
II-D Similarity-Dissimilarity Loss
To address the aforementioned challenges, we introduce the concepts of similarity and dissimilarity based on set-theoretic relations: (i) As depicted in Figure 1, Similarity represents the intersection of sets (i.e., $\mathcal{S}\cap\mathcal{T}$), and (ii) we define Dissimilarity as the set difference between $\mathcal{T}$ and the intersection $\mathcal{S}\cap\mathcal{T}$ with respect to sample $p$ (i.e., $\mathcal{T}-\mathcal{S}\cap\mathcal{T}$). For each anchor $i$, we formulate the Similarity-Dissimilarity Loss as:
$\mathcal{L}_{i}^{our}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (13)
@@ II-F Theoretical Analysis
Theorem 2.
$\mathcal{K}_{m}^{s}\mathcal{K}_{m}^{d}=1,\quad\text{when }m=2$ (21)
@@
Theorem 4.
$|\tilde{\boldsymbol{y}}_{p_{3}}|=|\tilde{\boldsymbol{y}}_{p_{4}}|$ (27)
$\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ (28)
@@
Theorem 5.
$|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|\leq|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$ (29)
$\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ (30)

### v4

@@ Abstract
Supervised contrastive learning has achieved remarkable success by leveraging label information; however, determining positive samples in multi-label scenarios remains a critical challenge. In multi-label supervised contrastive learning (MSCL), multi-label relations are not yet fully defined, leading to ambiguity in identifying positive samples and formulating contrastive loss functions to construct the representation space. To address these challenges, we: (i) first define five distinct multi-label relations in MSCL to systematically identify positive samples, (ii) introduce a novel Similarity-Dissimilarity Loss that dynamically re-weights samples through computing the similarity and dissimilarity factors between positive samples and given anchors based on multi-label relations, and (iii) further provide theoretical grounded proofs for our method through rigorous mathematical analysis that supports the formulation and effectiveness of the proposed loss function. We conduct the experiments across both image and text modalities, and extend the evaluation to medical domain. The results demonstrate that our method consistently outperforms baselines in a comprehensive evaluation, confirming its effectiveness and robustness. Code is available at: https://github.com/guangminghuang/similarity-dissimilarity-loss.
1 Introduction
@@
Supervised contrastive learning effectively utilizes label information to yield promising results in single-label scenarios [1, 2, 3]. However, identifying positive samples in multi-label supervised contrastive learning (MSCL) remains a challenge [4]. For example, consider a set of images containing cats and puppies, wherein an anchor image depicts a cat; in the single-label paradigm, positive and negative instances can be unambiguously delineated based on their corresponding taxonomic annotations. Conversely, MSCL introduces inherent classification ambiguity when determining whether an image containing both cats and puppies should be designated as a positive or negative sample in relation to the anchor.
@@
A critical question arises: Should a sample be considered positive when its label set partially overlaps with or exactly matches that of the anchor? Currently, three principal strategies exist for identifying positive samples in multi-label scenarios [4]: (i) ALL considers only samples with an exactly matching label set as positive; (ii) ANY identifies samples with any overlapping class with the anchor as positive, and (iii) MulSupCon [4] conceptually aligns with the ANY approach but treats each label independently, thereby generating multiple distinct positive sets for individual anchors.
@@
However, these methods have inherent limitations, since previous research has overlooked the complicated multi-label relations among samples in MSCL. As illustrated in Figure 1, we introduce five distinct set relations among samples to facilitate a more comprehensive identification of positive sets. The ALL strategy exclusively considers relation $R2$ while disregarding the potential contributions of $R3$, $R4$ and $R5$. Furthermore, long-tailed distributions, when tail samples serve as anchors, the ALL strategy’s requirement for exact label matches significantly impedes these tail anchors from identifying adequate positive samples within a limited batch size, potentially degenerating the method to unsupervised contrastive learning in extreme scenarios [5, 6, 7]. Conversely, both ANY and MulSupCon approaches treat relations $R2$, $R3$, $R4$, and $R5$ identically with equivalent weights in contrastive loss functions, which constitutes a suboptimal approach given the inherent differences among these relations. A detailed mathematical analysis of the limitations for these methods is presented in Section 2 and additional discussion of related work in Appendix F.
@@
Figure 1: Five distinct multi-label relations between samples and a given anchor. $\Omega$ denotes a universe that contains all label entities. Here is an example with five different relations between sample $p$ and anchor $i$, where the labels are represented as one-hot vectors.
@@
2 Methods
2.1 Preliminaries
Given a batch of $N$ randomly sample/label pairs, $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1,\dots,N}$, where $\boldsymbol{x}_{i}$ denotes the $i$-th sample and $\boldsymbol{y}_{i}$ its corresponding labels. Here, $\boldsymbol{y}_{i}=\{y_{i}^{(l)}\}_{l=1,\dots,L}$ represents the multi-labels of sample $i$, where $y_{i}^{(l)}$ denotes the $l$-th label of sample $i$ and $L$ is the total number of labels for sample $i$. After data augmentation, the training batch consists of $2N$ pairs, $\{\tilde{\boldsymbol{x}}_{j},\tilde{\boldsymbol{y}}_{j}\}_{j=1,\dots,2N}$, where $\tilde{\boldsymbol{x}}_{2i}$ and $\tilde{\boldsymbol{x}}_{2i-1}$ are two random augmentations of $\boldsymbol{x}_{i}$ ($i=1,\dots,N$) and $\tilde{\boldsymbol{y}}_{2i-1}=\tilde{\boldsymbol{y}}_{2i}=\boldsymbol{y}_{i}$. For brevity, we refer to this collection of $2N$ augmented samples as a "batch" [1].
2.2 Multi-label Supervised Contrastive Loss
@@
Zhang and Wu [4] propose an approach that considers each label $\tilde{y}_{i}^{(l)}$ independently, forming multiple positive sets for a given anchor sample $i$. For each label $\tilde{y}_{i}^{(l)}\in\tilde{\boldsymbol{y}}_{i}$, the positive set for the MulSupCon is defined as:
@@
$\mathcal{L}_{i}^{\text{mul}}=\sum_{\tilde{y}_{p}^{(l)}\in\tilde{\boldsymbol{y}}_{i}}\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (6)
@@
$\mathcal{L}^{\text{mul}}=\frac{1}{\sum_{i}|\tilde{\boldsymbol{y}}_{i}|}\sum_{i\in I}\mathcal{L}_{i}^{\text{mul}}$ (7)
2.3 Multi-label Relations
@@
For this example (see Figure 1), the MulSupCon successfully discriminates $R3$ and $R4$ from $R2$ and $R5$; however, it fails to establish a distinction between $R2$ and $R5$. This limitation arises primarily because MulSupCon exclusively considers the overlapping regions (Similarity  [^3] The definition of Similarity is introduced in Section 2.4) between anchor $i$ and sample $p$ (i.e., The intersection of sets $\mathcal{S}$ and $\mathcal{T}$), while disregarding the complementary non-intersecting domains (Dissimilarity  [^4] The definition of Dissimilarity is introduced in Section 2.4). That is to say, the similarity between positive samples and anchors is considers, but not yet dissimilarity, which is one of critical information for representation learning in MSCL.
@@
2.4 Similarity-Dissimilarity Loss
To address the aforementioned challenges, we introduce the concepts of similarity and dissimilarity based on set-theoretic relations: (i) As depicted in Figure 1, Similarity represents the intersection of sets (i.e., $\mathcal{S}\cap\mathcal{T}$), and (ii) we define Dissimilarity as the set difference between $\mathcal{T}$ and the intersection $\mathcal{S}\cap\mathcal{T}$ with respect to sample $p$ (i.e., $\mathcal{T}\setminus\mathcal{S}\cap\mathcal{T}$). For each anchor $i$, we formulate the Similarity-Dissimilarity Loss as:
$\mathcal{L}_{i}^{\text{our}}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{a}/\tau)}$ (13)
@@ 2.6 Theoretical Analysis
See Appendix A.1. ∎
Theorem 2.
$\mathcal{K}_{m}^{s}\mathcal{K}_{m}^{d}=1,\quad\text{when }m=2$ (20)
Proof.
See Appendix A.2. ∎
Theorem 3.
$10<\mathcal{K}_{m}^{s}\mathcal{K}_{m}^{d}<$ 1 (21)
Proof.
See Appendix A.3. ∎
Theorem 4.
$|\tilde{\boldsymbol{y}}_{p_{3}}|=|\tilde{\boldsymbol{y}}_{p_{4}}|$ (22)
$\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ (23)
Proof.
See Appendix A.4. ∎
Theorem 5.
$|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|\leq|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$ (24)
$\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ (25)
Proof.
See Appendix A.5. ∎
3 Experiments
We conduct the experiments to compare Similarity-Dissimilarity Loss with current contrastive loss functions (ALL, ANY, and MulSupCon) in a comprehensive evaluation, considering: (i) Data modality: image and text data; (ii) Domain-specific: general text data (AAPD) and medical domain (MIMIC III and IV); (iii) Data distribution: full setting (extreme long-tailed distribution) and top-50 frequent labels setting; (iv) ICD code versions: ICD-9 and ICD-10, and (v) Models: ResNet-50, RoBERTa-based, Llama-3.1-8B, and PLM-ICD. For comprehensive information, please refer to Appendix B for detailed dataset descriptions and evaluation metrics, Appendix C for baseline models and encoder architectures, and Appendix D for implementation details.
4 Results and Analysis
More experiment results for the evaluation on medical domain in Appendix E
4.1 Evaluation on Image
The experimental results in Table 1 demonstrate that our proposed loss function outperforms baselines across all metrics, including micro-F1, macro-F1, and mAP, on all image datasets (MS-COCO, PASCAL, and NUS-WIDE). Compared to MulSulCon, our method achieves significant improvements of $2.07/3.78/1.51$ in micro-/macro-F1/mAP on MS-COCO and $1.47/3.13/4.27$ on NUS-WIDE.
Figure 2(a) illustrates the comparison between Similarity-Dissimilarity Loss and MulSupCon as measured by micro- and macro-F1 metrics. The results indicate that our method yields substantially greater improvements in macro-F1 compared to micro-F1 across all image datasets. Specifically, macro-F1 increases by $5.7\%$ on MS-COCO and $5.8\%$ on NUS-WIDE, whereas micro-F1 exhibits more modest improvements of $2.9\%$ and $2.0\%$, respectively. Macro-F1 assigns equal importance to each class regardless of its frequency, rendering it particularly appropriate for evaluating performance on imbalanced datasets where minority class prediction accuracy is critical [8, 7]. In contrast, micro-F1 places more considerable weight on classes with more samples, making it more appropriate when larger classes should have a more potent influence on the overall score [8, 9]. Multi-label classification inherently faces more pronounced challenges with long-tailed distributions than single-label classification due to exponential output space complexity, intricate label co-occurrence patterns, and high annotation costs [7]. The observed superior improvement in macro-F1 metrics provides compelling evidence that our method demonstrates exceptional efficacy in addressing long-tailed distribution challenges, a capability particularly crucial in multi-label scenarios.
However, on the PASCAL dataset, our method demonstrates mere marginal improvements, with gains of $0.88/0.84/0.17$ in micro/macro-F1/mAP, respectively. This limited enhancement can be attributed to the structural characteristics of PASCAL, wherein the average number of labels per instance is approximately $1.5$ (as detailed in Table 3), causing the task to approximate single-label classification, particularly when the batch size is limited [1]. Consequently, loss functions specifically designed for multi-label scenarios exert minimal influence on model performance under these conditions. As Audibert et al. [10] have demonstrated, the cardinality of the label space constitutes a significant determinant of model efficacy within MSCL .
Furthermore, Figure 2(b) reveals that the standard deviation across four methods for PASCAL equals $0.58/0.64/0.27$ in micro/macro-F1/mAP, which are considerably lower than the corresponding standard deviations observed for the MS-COCO and NUS-WIDE. This statistical finding suggests that the efficacy of specialized multi-label loss functions diminishes significantly when the average label cardinality per instance approaches $1$ in MSCL. This finding further corroborates our theoretical analysis and hypothesis in the Section 2, wherein Similarity-Dissimilarity Loss degenerates to single-label scenarios (see Eq. (17)).
Table 1: Results on image datasets. We compare our method with baselines (ALL, ANY and MulSupCon). $\mu$-F1 and $\nu$-F1 present as micro-F1 and macro-F1, respectively.
Method MS-COCO PASCAL NUS-WIDE
$\mu$-F1 $\nu$-F1 mAP $\mu$-F1 $\nu$-F1 mAP $\mu$-F1 $\nu$-F1 mAP
ALL 68.93 63.32 64.11 82.53 79.87 79.32 70.25 52.84 51.35
ANY 64.80 57.37 56.90 82.31 79.65 79.15 68.42 50.65 49.28
MulSupCon 71.33 66.25 67.69 82.75 80.26 79.58 71.88 54.36 52.47
Ours 73.40 70.03 69.20 83.63 81.10 79.75 73.35 57.49 56.74
Figure 2: (a) Comparison of performance improvements between Similarity-Dissimilarity Loss and MulSupCon with micro- and macro-F1 metrics. (b) Comparison standard deviation of image datasets on micro-F1, macro-F1 and mAP metrics.
4.2 Evaluation on Text
We further evaluate our method on general text data, and the results demonstrate that our proposed loss function consistently surpasses baseline methods for both RoBERTa and Llama models across all metrics on the AAPD dataset (See Table 2). In contrast to the significant performance gains observed on image data, Similarity-Dissimilarity Loss achieves more modest enhancements of $0.90/1.79$ in micro/macro-F1 scores on RoBERTa, and $0.89/1.84$ on Llama. This attenuated performance differential can be attributed to the extensive knowledge already encoded within LLMs through their comprehensive pre-training paradigms [11].
Table 2: Results on AAPD Dataset. We compare our proposed Similarity-Dissimilarity Loss with baselines on general text data using RoBERTa-based and Llama-3.1-8B models
Method RoBERTa Llama
Micro-F1 Macro-F1 Micro-F1 Macro-F1
ALL 73.23 59.41 74.32 60.47
ANY 72.31 58.55 73.41 59.63
MulSupCon 73.64 60.52 74.72 61.58
Ours 74.54 62.31 75.61 63.42
5 Conclusion
In this paper, we introduce multi-label relations and propose a Similarity-Dissimilarity Loss for MSCL, which dynamically re-weights the loss by the similarity-dissimilarity factor. We provide theoretical grounded proof for our method through rigorous mathematical analysis that supports the formulation and effectiveness of the proposed loss function. The experimental results further confirm the effectiveness and robustness of our method in image, text and medical domain. Despite the advancements, limitations warrant acknowledgment. The Similarity-Dissimilarity Loss, while effective via linear factors in MSCL, may benefit from further explore non-linear similarity-dissimilarity weighting factors and learnable hyperparameter for contrastive loss function to capture more complex semantic relations and long-tailed distribution.
Appendix A Theorem Proofs
A.1 Proof of Theorem 1
See 1
Proof.
@@ A.1 Proof of Theorem 1
A.2 Proof of Theorem 2
See 2
@@ A.3 Proof of Theorem 3
A.4 Proof of Theorem 4
See 4
@@ A.4 Proof of Theorem 4
A.5 Proof of Theorem 5
See 5